# Progressive Mixed-Precision Decoding for Efficient LLM Inference

### v1

@@ Abstract
In spite of the great potential of large language models (LLMs) across various tasks, their deployment on resource-constrained devices remains challenging due to their excessive computational and memory demands. Quantization has emerged as an effective solution by storing weights in reduced precision. However, utilizing low precisions (i.e. 2/3-bit) to substantially alleviate the memory-boundedness of LLM decoding, still suffers from prohibitive performance drop. In this work, we argue that existing approaches fail to explore the diversity in computational patterns, redundancy, and sensitivity to approximations of the different phases of LLM inference, resorting to a uniform quantization policy throughout. Instead, we propose a novel phase-aware method that selectively allocates precision during different phases of LLM inference, achieving both strong context extraction during prefill and efficient memory bandwidth utilization during decoding. To further address the memory-boundedness of the decoding phase, we introduce Progressive Mixed-Precision Decoding (PMPD), a technique that enables the gradual lowering of precision deeper in the generated sequence, together with a spectrum of precision-switching schedulers that dynamically drive the precision-lowering decisions in either task-adaptive or prompt-adaptive manner. Extensive evaluation across diverse language tasks shows that when targeting Nvidia GPUs, PMPD achieves 1.4--12.2$\times$ speedup in LLM linear layers over fp16 models, while when targeting an LLM-optimized NPU, our approach delivers a throughput gain of 3.8--8.0$\times$ over fp16 models and up to 1.54$\times$ over uniform quantization approaches while preserving the output quality.
@@ 1 Introduction
To enhance the hardware performance of LLM inference at the edge, several approaches have been proposed, including model compression and adaptive inference techniques. Particularly, quantization has been widely studied, to adopt low-precision representations for the weights (Yu et al., 2021) and/or activations (Tan et al., 2024), as well as its Key-Value (KV) cache (Hooper et al., 2024). Nevertheless, most existing approaches apply a uniform quantization scheme across both prefill and decoding stages, which presents two key limitations: i) They overlook the distinct fault tolerances against quantization errors in prefill and decoding stages, leading to a significant accuracy drop while reducing the bit-width for mobile-scale LLMs; ii) Since the decoding phase is memory-bound, reducing the bitwidth in this stage is more effective in improving hardware performance. As such, previous uniform quantization approaches can only achieve limited speedups, underscoring the need for more targeted quantization strategies specifically tailored to LLM decoding.
@@ 3.1 Observation One: Different Error Resilience in Prefill and Decoding
Figure 2: Phase-aware Precision Allocation. Here prefilling was performed by 2- (top) or 3-bit (bottom) Vicuna-7B, while decoding employed the 2-bit model for both cases. We observe three key improvements. i) enhanced reasoning abilities. The model demonstrated higher accuracy for numerical tasks, such as calculating Fibonacci numbers (left). ii) improved instruction-following. The model understood and responded to user instructions (middle). iii) strengthened emerging abilities like creativity. The model successfully handled open-ended tasks such as poem writing (right).
@@ 4.1 Phase-Aware Precision Allocation
The proposed Phase-Aware Precision Allocation capitalizes on the unique characteristics of each LLM inference phase and tailors accordingly the per-phase weight precision. Specifically, our inference scheme dictates that a high-precision model is used during the prefill phase – allowing the extraction of rich context from the input prompt – which in turn enables the use of lower-precision weights in the memory-bound decoding phase, and thus yields more efficient utilization of the memory bandwidth. As discussed in Section 3, since the prefill phase is compute-bound and the LLM inference latency is typically dominated by the decoding phase, using a high-precision model during prefill introduces negligible latency overhead.
@@ 4.2 Progressive Mixed-Precision Decoding
Input: LLM $m$ with max context length $CL$
@@ 23      $p^{\text{new}}\leftarrow p^{\text{decode}}$
24      for $i\leftarrow 0$ to $CL-1$ do $\triangleright$ Decoding Phase
@@ 29            $p^{\text{new}}\leftarrow\text{PMPDScheduler}(i+1)$triangleright Precision-switching scheduler
We parametrize PMPD with i) the precision set $\mathcal{P}$ and ii) the precision-switching schedule $S=\{st(p)\leavevmode\nobreak\ |\leavevmode\nobreak\ p\in\mathcal{P}\}$, which consists of the possible switching points of each precision defined as $st(p)\in[1,CL)$, where $CL$ is the maximum context length of a given LLM. Here, the starting point $st(p)$ is defined as the token index of the output sequence where we perform a switch from the current precision to the lower precision $p$. Under this setup, to obtain the highest-performing PMPD configuration, we pose the following optimization problem:
@@ 4.3 Precision-Switching Scheduler
Given the set of variably quantized models $\{m_{p}\}$ that was generated in the phase-aware precision allocation step (Section 4.1 and line 1), the objective function of Eq. (1) can be evaluated for all combinations of precision-switching schedules. Ideally, the highest-performing schedule considers all $p\in\mathcal{P}$ per-prompt per-step when deciding for a precision switch without the precedence constraint shown in Eq. (3). This schedule can be obtained through an exhaustive enumeration over ${|\mathcal{P}|}^{CL}$ schedules. However, for realistic values of $|\mathcal{P}|$ and $CL$, this computation quickly becomes intractable. Besides, even if such a schedule can be efficiently found, it is, in most cases, impractical as it would require excessively frequent switching, e.g. per decoding step, between variably quantized weights, aggravating the memory bandwidth demand (Section 3.2).
@@
Instead, through our approach of progressive precision lowering during decoding, the total number of schedules to be examined is reduced to $\sum^{|\mathcal{P}|-1}_{r=0}\frac{CL!}{r!(CL-r)!}\frac{(|\mathcal{P}|-1)!}{r!(|\mathcal{P}|-1-r)!}$ where $r$ is the number of times the precision switches. Nonetheless, the search time for the optimal schedule increases with $CL$ and $|\mathcal{P}|$, which is still prohibitive at run time. Hence, to further reduce the search time for the optimal $S$ to $O(1)$, we parameterize the number of candidate precision-switching points, denoted by $N$, and add a range constraint to the second constraint in Eq. (2) as $st(p)\in\{0\}\cup\{\frac{CL}{(N-1)},\frac{2CL}{(N-1)},...,\frac{(N-2)CL}{(N-1)},CL\}$. Under this setup, we propose two types of schedulers: i) a prompt-agnostic static scheduler, where $S$ is determined prior to deployment, and ii) a task-agnostic learned scheduler, where $S$ is derived dynamically for each prompt.
@@
Hence, we utilize the features extracted in the prefill phase as inputs to our scheduler. With this approach, we are able to maintain a lightweight architecture for the scheduler and introduce minimal overhead during inference as the scheduler generates schedule $S$ once before decoding (line 13). In Section 5.3, we investigate different input features, i.e. KV caches and activations, from various layers of the model and find that utilizing the prefilled KV cache is efficacious in most cases, supporting our claim in Section 3.1 that the prefilled KV cache provides good task-specific context. We also find that despite not requiring a task-specific validation set and labeled ground truths, our learned scheduler performs close to the static and, in some cases, even outperforms it in both algorithmic and hardware performance. Its architecture and training are detailed in Appendix A.1.
@@ 5 Evaluation
Models and Datasets. We conducted experiments on edge-deployable models, including Vicuna-7B (Chiang et al., 2023), MobileLLaMA-1.4B (Chu et al., 2023), Stable LM Zephyr-3B [^1] https://stability.ai/news/stablelm-zephyr-3b-stability-llm, and Phi-1.5 (Li et al., 2023), evaluating their zero-shot generative performance on news summarization, dialogue summarization, and translation tasks using the CNN/DM (Hermann et al., 2015), Dialogsum (Chen et al., 2021), and IWSLT French-English datasets (Cettolo et al., 2017), respectively. These tasks encompass context understanding and text generation, providing a comprehensive evaluation of PMPD [^2] Hereafter, PMPD refers to our approach including both phase-aware precision allocation and progressive mixed-precision decoding.. We also tested open-ended question answering on MT-Bench (Chiang et al., 2023), a special case where Static scheduler is infeasible due to the absence of a validation set.
@@ 5.1 Algorithmic Performance
Table 1: Performance comparison of Static and Learned schedulers against low-precision (baseline-l) and high-precision (baseline-h) baselines, as well as DNS with the same average bitwidth as the scheduler with highest performance. For CNN/DM and Dialogsum, models used from top to bottom are Vicuna-7B (2/3 bit), MobileLLaMA (3/4 bit), and Phi-1.5 (2/4 bit). For IWSLT, the models are Vicuna-7B, MobileLLaMA, and Zephyr-3B (3/4 bit). Pairwise winners between the scheduler and DNS are underlined, while the highest overall scores are in bold.
@@ A.1.1 Model Architecture
To avoid runtime overhead, we employ a lightweight attention module for Learned scheduler. Given an input Keys cache $K\in\mathbb{R}^{T\times D_{k}}$ and Values cache $V\in\mathbb{R}^{T\times D_{v}}$, the output of Learned scheduler is computed as:
@@
where $MLP$ is a feedforward neural network with one hidden layer and a ReLU as its activation, and $q_{\theta}\in\mathbb{R}^{D_{v}}$ is a learned query vector. In other words, the trainable parameters of the network include the query vector and the parameters of the MLP.
@@ A.1.2 Training Approach
To prepare the training dataset, we use the C4 dataset as the seed dataset and randomly truncate each sample to create a completion task. For each sample, we generate $N$ sequences, each corresponding to a high-precision step from $\{0,\frac{CL}{N-1},\frac{2CL}{N-1},\ldots,\frac{(N-2)CL}{N-1},CL\}$. The ground-truth label is set to the lowest high-precision step such that the Rouge-L score of the sequence matches or exceeds that of the fp16 model output.
@@
A.2 GPU Matrix Multiplication Latency Evaluation

### v2

@@ Abstract
Despite the great potential of large language models (LLMs), deploying them on resource-constrained devices is challenging due to high computational and memory demands. Quantization has emerged as an effective solution by storing weights in lower precision. However, Post-Training Quantization (PTQ) at extremely low precisions (i.e., 2/3-bit) still incurs severe performance degradation. In this work, we argue that existing approaches fail to explore the diversity in computational patterns, redundancy, and sensitivity to approximations of the different phases of LLM inference, resorting to a uniform quantization policy throughout. Instead, we propose a novel phase-aware method that selectively allocates precision during different phases of LLM inference, achieving both strong context extraction during prefill and efficient memory bandwidth utilization during decoding. To further address the memory-boundedness of the decoding phase, we introduce Progressive Mixed-Precision Decoding (PMPD), a technique that enables the gradual lowering of precision deeper in the generated sequence, together with a spectrum of precision-switching schedulers that dynamically drive the precision-lowering decisions in either task-adaptive or prompt-adaptive manner. Extensive evaluation across diverse language tasks shows that when targeting Nvidia GPUs, PMPD achieves 1.4--12.2$\times$ speedup in LLM linear layers over fp16 models and up to 1.41$\times$ over uniform quantization. When targeting an LLM-optimized NPU, our approach delivers a throughput gain of 3.8--8.0$\times$ over fp16 models and up to 1.54$\times$ over uniform quantization approaches while preserving the output quality. Our code is available at github.com/SamsungLabs/PMPD.
@@ 1 Introduction
To enhance the hardware performance of LLM inference at the edge, techniques like model compression and adaptive inference have been explored. Quantization, in particular, reduces precision for weights (Yu et al., 2021), activations (Tan et al., 2024), and KV cache (Hooper et al., 2024). However, most methods apply uniform quantization across prefill and decoding, leading to two key issues: i) Ignoring distinctive fault tolerances in prefill and decoding stages, causing accuracy drops while reducing bit-widths for mobile-scale LLMs; ii) Decoding is memory-bound, so bit-width reduction is more effective here for speedup. Thus, uniform quantization yields limited gains, highlighting the need for more targeted quantization strategies specifically tailored to LLM decoding.
@@ 3.1 Observation One: Different Error Resilience in Prefill and Decoding
Figure 2: Phase-aware Precision Allocation. Here prefilling was performed by 2- (top) or 3-bit (bottom) Vicuna-7B, while decoding employed the 2-bit model for both cases. We observe three key improvements. i) enhanced reasoning abilities. The model demonstrated higher accuracy for numerical tasks, such as calculating Fibonacci numbers (left). ii) improved instruction-following. The model understood and responded to user instructions (middle). iii) strengthened emerging abilities like creativity. The model successfully handled open-ended tasks such as poem writing (right). More quantitative results are shown in Appendix A.2 and Figure 6.
@@ 4.1 Phase-Aware Precision Allocation
The proposed Phase-Aware Precision Allocation capitalizes on the unique characteristics of each LLM inference phase and tailors accordingly the per-phase weight precision. Specifically, our inference scheme dictates that a high-precision model is used during the prefill phase which in turn enables the use of lower-precision weights in the memory-bound decoding phase, and thus yields more efficient utilization of the memory bandwidth. Since the prefill phase is compute-bound and the LLM inference latency is typically dominated by the decoding phase, using a high-precision model during prefill introduces negligible latency overhead.
@@ 4.2 Progressive Mixed-Precision Decoding
Input: LLM $m$ with max output length $OL$
@@ 23      $p^{\text{new}}\leftarrow p^{\text{decode}}$
24      for $i\leftarrow 0$ to $OL-1$ do $\triangleright$ Decoding Phase
@@ 29            $p^{\text{new}}\leftarrow\text{PMPDScheduler}(i+1)$triangleright Precision-switching scheduler
We parametrize PMPD with i) the precision set $\mathcal{P}$ and ii) the precision-switching schedule $S=\{st(p)\leavevmode\nobreak\ |\leavevmode\nobreak\ p\in\mathcal{P}\}$, which consists of the possible switching points of each precision defined as $st(p)\in[1,OL)$, where $OL$ is the maximum output length of a given LLM. Here, the starting point $st(p)$ is defined as the token index of the output sequence where we perform a switch from the current precision to the lower precision $p$. Under this setup, to obtain the highest-performing PMPD configuration, we pose the following optimization problem:
@@ 4.3 Precision-Switching Scheduler
Given the set of variably quantized models $\{m_{p}\}$ that was generated in the phase-aware precision allocation step (Section 4.1 and line 1), the objective function of Eq. (1) can be evaluated for all combinations of precision-switching schedules. Ideally, the highest-performing schedule considers all $p\in\mathcal{P}$ per-prompt per-step when deciding for a precision switch without the precedence constraint shown in Eq. (3). This schedule can be obtained through an exhaustive enumeration over ${|\mathcal{P}|}^{OL}$ schedules. However, for realistic values of $|\mathcal{P}|$ and $OL$, this computation quickly becomes intractable. Besides, even if such a schedule can be efficiently found, it is, in most cases, impractical as it would require excessively frequent switching, e.g. per decoding step, between variably quantized weights, aggravating the memory bandwidth demand (Section 3.2).
@@
Instead, through our approach of progressive precision lowering during decoding, the total number of schedules to be examined is reduced to $\sum^{|\mathcal{P}|-1}_{r=0}\frac{OL!}{r!(OL-r)!}\frac{(|\mathcal{P}|-1)!}{r!(|\mathcal{P}|-1-r)!}$ where $r$ is the number of times the precision switches. Nonetheless, the search time for the optimal schedule increases with $OL$ and $|\mathcal{P}|$, which is still prohibitive at run time. Hence, to further reduce the search time for the optimal $S$ to $O(1)$, we parameterize the number of candidate precision-switching points, denoted by $N$, and add a range constraint to the second constraint in Eq. (2) as $st(p)\in\{0\}\cup\{\frac{OL}{(N-1)},\frac{2OL}{(N-1)},...,\frac{(N-2)OL}{(N-1)},OL\}$. Under this setup, we propose two types of schedulers: i) a prompt-agnostic static scheduler, where $S$ is determined prior to deployment, and ii) a task-agnostic learned scheduler, where $S$ is derived dynamically for each prompt.
@@
Hence, we utilize the features extracted in the prefill phase as inputs to our scheduler. With this approach, we are able to maintain a lightweight architecture for the scheduler and introduce minimal overhead during inference as the scheduler generates schedule $S$ once before decoding (line 13). In Section 5.3, we investigate different input features, i.e. KV caches and activations, from various layers of the model and find that utilizing the prefilled KV cache is efficacious in most cases. Given an input Key cache $K\in\mathbb{R}^{T\times D_{k}}$ and Value cache $V\in\mathbb{R}^{T\times D_{v}}$, the output of Learned scheduler is computed as $O=\text{MLP}(\text{Softmax}(\frac{q_{\theta}K^{T}}{\sqrt{D_{k}}})V)$, where $q_{\theta}\in\mathbb{R}^{D_{k}}$ is a learned query vector.
We also find that despite not requiring a task-specific validation set and labeled ground truths, our learned scheduler performs close to the static and, in some cases, even outperforms it in both algorithmic and hardware performance. Its architecture and training are detailed in Appendix A.1.
@@ 5 Evaluation
Models and Datasets. We conducted experiments on edge-deployable models, including Vicuna-7B (Chiang et al., 2023), MobileLLaMA-1.4B (Chu et al., 2023), Stable LM Zephyr-3B [^1] https://stability.ai/news/stablelm-zephyr-3b-stability-llm, and Phi-1.5 (Li et al., 2023b), evaluating their zero-shot generative performance on news summarization, dialogue summarization, and translation tasks using the CNN/DM (Hermann et al., 2015), Dialogsum (Chen et al., 2021), and IWSLT French-English datasets (Cettolo et al., 2017), respectively. These tasks encompass context understanding and text generation, providing a comprehensive evaluation of PMPD [^2] Hereafter, PMPD refers to our approach including both phase-aware precision allocation and progressive mixed-precision decoding.. We also tested open-ended question answering on MT-Bench (Chiang et al., 2023), a special case where Static scheduler is infeasible due to the absence of a validation set.
@@ 5.1 Algorithmic Performance
Table 1: Performance comparison of Static and Learned schedulers against low-precision (baseline-l) and high-precision (baseline-h) baselines, as well as DNS with the same average bitwidth as the scheduler with highest performance. For CNN/DM and Dialogsum, models used from top to bottom are Vicuna-7B (2/3 bit), MobileLLaMA (3/4 bit), and Phi-1.5 (3/4 bit). For IWSLT, the models are Vicuna-7B, MobileLLaMA, and Zephyr-3B (3/4 bit). Pairwise winners between the scheduler and DNS are underlined, while the highest overall scores are in bold.
@@ A.1.1 Model Architecture
To avoid runtime overhead, we employ a lightweight attention module for Learned scheduler. Given an input Key cache $K\in\mathbb{R}^{T\times D_{k}}$ and Value cache $V\in\mathbb{R}^{T\times D_{v}}$, the output of Learned scheduler is computed as:
@@
where $MLP$ is a feedforward neural network with one hidden layer and a ReLU as its activation, and $q_{\theta}\in\mathbb{R}^{D_{k}}$ is a learned query vector. In other words, the trainable parameters of the network include the query vector and the parameters of the MLP.
@@ A.1.2 Training Approach
To prepare the training dataset, we use the C4 dataset as the seed dataset and randomly truncate each sample to create a completion task. For each sample, we generate $N$ sequences, each corresponding to a high-precision step from $\{0,\frac{OL}{N-1},\frac{2OL}{N-1},\ldots,\frac{(N-2)OL}{N-1},OL\}$. The ground-truth label is set to the lowest high-precision step such that the Rouge-L score of the sequence matches or exceeds that of the fp16 model output.
@@
A.2 Error Resilience of Prefill and Decode Bitwidth
Decoding/Prefill Strategy Rouge-L
2bit Prefill, 2bit Decode 18.3
3bit Prefill, 2bit Decode 28.1
2bit Prefill, 3bit Decode 22.2
Table 4: Rouge-L scores for different decoding and prefill bitwidth allocations. Vicuna-7B and MT-Bench are used. Outputs from fp16 model are used as reference answer.
As shown in Table 4, we observed that using a higher bitwidth for prefilling significantly improves output quality. On the other hand, using higher bitwidth for decoding does not improve the performance as much. This demonstrates that a higher bitwidth should be used for prefilling.
A.3 Performance of PMPD for Larger Models
Method CNN/DM Dialogsum IWSLT
Model ($\downarrow$): Llama3-8B Model ($\downarrow$)): Llama3-8B Model ($\downarrow$)): Llama3-8B
Longchat-13B Longchat-13B Longchat-13B
Bit Rouge-L/ Bit Rouge-L/ Bit BLEU/
BERTScore BERTScore SacreBLEU
Baseline-l 2 4.35 / 75.8 2 5.58 / 81.7 2 0 / 0.16
Baseline-h 3 18.3 / 85.2 3 19.0 / 85.8 3 23.1 / 23.2
PMPD-Static 2.66 17.8 / 84.8 2.31 18.7 / 85.8 2.71 22.1 / 22.1
Baseline-l 2 9.29 / 72.8 2 10.6 / 80.8 2 2.94 / 2.94
Baseline-h 3 19.9 / 86.4 3 19.9 / 86.6 3 21.6 / 21.6
PMPD-Static 2.37 19.9 / 86.4 2.0 20.3 / 86.5 2.71 21.4 / 21.4
Table 5: Performance of PMPD on models larger than 7B.
While PMPD mainly targets edge and mobile applications, it would be interesting to see if larger models could also benefit from PMPD. We conducted additional experiments using Llama3-8B (AI@Meta, 2024) and longchat-16k-13B (Li et al., 2023a) models, with a static scheduler for PMPD given the availability of validation data. The results shown in Table 5 demonstrate our method’s generalizability across model scales.
A.4 GPU End-to-End Latency Speedup
GPU Method Vicuna-7B MobileLlama Phi-1.5 Zephy-3B
4090 Baseline-h 3.05×3.05$\times$3.05 × 1.80×1.80$\times$1.80 × 1.75×1.75$\times$1.75 × 2.07×2.07$\times$2.07 ×
PMPD 3.36×3.36$\times$3.36 × 1.90×1.90$\times$1.90 × 1.83×1.83$\times$1.83 × $2\times2$
A40 Baseline-h 2.67×2.67$\times$2.67 × 1.71×1.71$\times$1.71 × 1.66×1.$6$ 1.91×1.91$\times$1.91 ×
PMPD 2.91×2.91$\times$2.91 × 1.82×1.82$\times$1.82 × 1.74×1.74$\times$1.74 × 2.07×2.07$\times$2.07 ×
Table 6: End-to-End GPU latency speedup as compared to FP16 model.
We provide the GPU end-to-end latency speedup as compared to FP16 models in Table 6. PMPD consistently speedup improvement over the uniform quantization baseline upto 0.31 $\times$.
A.5 GPU Matrix Multiplication Latency Evaluation
@@ A.7 Alternative PTQ Method
A.8 Example Outputs by PMPD
In this section, we include some example outputs by Vicuna-7B with PMPD on MT-Bench, The precisions used are 3 and 4 bits.
A Success Case
Question: How do the stages of life shape our understanding of time and mortality?
Answer: The stages of life can shape our understanding of time and mortality in several ways.
1. Childhood: Children are often more focused on the present moment and less concerned with the future. This can lead to a more relaxed and flexible understanding of time and mortality.
2. Adulthood: As we enter into adulthood, we become more aware of the passage of time and the possibility of death. This can lead to a more pessimistic view of time and mortality.
3. Midlife: Midlife is a time of significant life changes, including career changes, family responsibilities, and the possibility of death. This can lead to a more realistic and optimistic view of time and mortality.
4. Later life: As we age, we may experience changes in our lives and the possibility of death. This can lead to a more realistic and pessimistic view of time and mortality.
Overall, the stages of life can shape our understanding of time and mortality by shaping our understanding of the present moment and the possibility of death.
The PMPD model is able to produce sensible results for humanity subjects.
A Failure Case
Question: When a number is divided by 10, the remainder is 4. What is the remainder when twice the number is divided by 4?
Answer: To find the remainder when a number is divided by 4, you can use the following steps:
1. Write down the number you want to find the remainder of.
2. Write down the number you want to find the remainder of again.
3. Multiply the number you wrote down 2 times.
4. Write down the result.
Here, while coherent texts are generated, the answer is not very relevant to the question.